INAF queue observing, 2013 Nov 01-Nov 02

Observer: M. Fumana, R. Speziali
Telescope Operator: G. Becchetti
Support Astronomer: M. Edwards / H. Khun (on call)
Instrument: MODS (LUCI)

Summary

At the sunset clear sky.
No MODS available at the beginning of the night.
Three primary mirror panic events during the night Problem in LUCI grism tilting. We had to skip a high priority program

Data Summary

Project_RUN Objects Filter ExpTime Notes
FusiPecci _07 07M Dual grating 45min  
FusiPecci _07 07N Dual grating 45min  
FusiPecci _07 07O Dual grating 45min  
FusiPecci _07 07P Dual grating 1h15min PROGRAM COMPLETED
Vanzella_20A macsj0717 Dual grating 2h20min  

Details

00:33 The TO opens the dome.

01:04 acquiring Ks sky flats 004-008

01:24 Primary mirror in panic! "AAAHAHAHAAHAHAGH!!"

01:49 starting acquisition Cresci 8A

02:03 Primary mirror in panic again! "AAAHAHAHAAHAHAGH!!" ma **^^#@@##@!!

02:18 starting acquisition Cresci 8A

02:33 executeLUCIscript

02:34 Error in the grating unit: error to tilt the grating

02:40 We are not able to stabilize the grating a 2.138

Problem with MODS software control solved. The weather forecast says it will snow and clouds will come during the next nights. We try to complete at least some MODS programs.

02:42 Switch to MODS

03:09 FusiPecci _07M: acqMODS acqMODS B27_ls.acq
  • Manual offset (-0.1,0)

03:22 execMODS B27_ls_12.obs
  • seeing ~1"
  • gcs seeing ~1.2"

04:15 FusiPecci _07N: acqMODS B232_ls_PA220.acq

04:27 execMODS MODS B232_ls_13.obs
  • seeing ~1.1

05:18 FusiPecci _07O: acqMODS B45_ls_PA150.acq
  • manual offset (-0.1,0)

05:32 execMODS B45_ls_14.obs
  • seeing ~1.1

06:24 FusiPecci _07P: acqMODS MCGC5_ls_PA93.acq

06:37 execMODS MCGC5_ls_15.obs
  • seeing ~1.3"
  • seeing is raising ~1.4"
  • very variable seeing peak of ~1.7"
  • in the 3rd exposure seeing is going down ~1.1"

07:36 Because of the seeing in the first exposures, we acquire again 2 frames updating PA in the script. We also try to use a brighter guiding star

07:37 FusiPecci _07P: acqMODS MCGC5_ls_PA93.acq

06:49 execMODS MCGC5_ls_15.obs
  • seeing ~ 1.1-1.2"
  • Stop after 2 exposures ==

08:26 We redo FusiPecci _07I done yesterday with a bad seeing

08:27 FusiPecci _07I: acqMODS B311_ls_PA84.acq

08:33 Left primary mirror has panic

09:16 TO solved problem on the mirror

09:16 Slew to Vanzella

09:22 acqMODS mods.macsj0717_mos.acq

09:41 execMODS mods.macsj0717_mos_3.obs
  • seeing ~1"

10:49 execMODS mods.macsj0717_mos_4.obs
  • seeing ~1.2"

11:58 Unwrap the rotator

12:00 acqMODS mods.macsj0717_mos.acq

12:21 execMODS mods.macsj0717_mos_5.obs
  • seeing ~1"
  • Acquisition aborted after 1 exposure, because of the twilight

12:55 execMODS modsSleep.pro

13:00 Dome close

-- MarcoFumana - 02 Nov 2013
Topic revision: r7 - 02 Mar 2014, SusannaBisogni
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback